he following information applies to the questions displayed below.]

In 2021, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2023. Information related to the contract is as follows:

2021 2022 2023
Cost incurred during the year $ 2,291,000 $ 3,555,000 $ 2,259,400
Estimated costs to complete as of year-end 5,609,000 2,054,000 0
Billings during the year 2,110,000 3,736,000 4,154,000
Cash collections during the year 1,855,000 3,500,000 4,645,000


Westgate recognizes revenue over time according to percentage of completion.

2-a. In the journal below, complete the necessary journal entries for the year 2021 (credit "Various accounts" for construction costs incurred).
2-b. In the journal below, complete the necessary journal entries for the year 2022 (credit "Various accounts" for construction costs incurred).
2-c. In the journal below, complete the necessary journal entries for the year 2023 (credit "Various accounts" for construction costs incurred).
Answer #1

Solution:

Workings are given at the end of the solution.

a. Journal entries for 2021:

2021
Debit Credit
Construction in progress 2,291,000
Various accounts 2,291,000
To record construction costs
Accounts receivable 2,110,000
Contract Billings 2,110,000
Recording Progress billings
Cash 1,855,000
Accounts receivable 1,855,000
Recording Collections
Work in Progress 609,000
(gross profit)
Construction Costs 2,291,000
(cost incurred)
Income (Revenue) from long-term contracts 2,900,000
Recording Gross Profit

b. Journal Entries for 2022:

2022
Debit Credit
Construction in progress 3,555,000
Various accounts 3,555,000
To record construction costs
Accounts receivable 3,736,000
Contract Billings 3,736,000
Recording Progress billings
Cash 3,500,000
Accounts receivable 3,500,000
Recording Collections
Work in Progress 945,000
(gross profit)
Construction Costs 3,555,000
(cost incurred)
Income (Revenue) from long-term contracts 4,500,000
Recording Gross Profit

c. Journal entries for 2023:

2023
Debit Credit
Construction in progress 2,259,400
Various accounts 2,259,400
To record construction costs
Accounts receivable 41,54,000
Contract Billings 41,54,000
Recording Progress billings
Cash 4,645,000
Accounts receivable 4,645,000
Recording Collections
Work in Progress 340,600
(gross profit)
Construction Costs 2,259,400
(cost incurred)
Income (Revenue) from long-term contracts 2,600,000
Recording Gross Profit

Workings:

Computation of Revenue:

    2021:       $2,291,000

                                        = 29.0% × $10,000,000 = $2,900,000

                     $7,900,000

    2022:       $5,846,000

                                        = 74.0% × $10,000,000 – 2,900,000 = $4,500,000

                     $7,900,000

    2023:     $10,000,000 – 7,400,000 = $2,600,000

Computation of Gross profit (loss):

     2021: $2,900,000 – 2,291,000 = $609,000

     2022: $4,500,000 – 3,555,000 = $945,000

     2023: $2,600,000 – 2,259,400 = $340,600